# SafeFix: Targeted Model Repair via Controlled Image Generation

### v1

@@ F.3 Evaluation with Mixed Train–Validation Synthetic Split
1,000 63.38 82.63
5,000 76.65 90.35
10,000 81.66 91.21
Table 6: Validation accuracy (%) on NN synthetic images, trained from scratch.
In our main experiments, we added $N$ validated synthetic images to $\mathcal{D}_{\mathrm{train}}$ while keeping $\mathcal{D}_{\mathrm{val}}$ fixed. To further examine the end-to-end value of our augmentation, we instead randomly split the $N$ synthetic images into an 80:20 train–val partition and trained ResNet from scratch. As shown in Table 6, our method consistently outperforms HiBug by large margins across all settings. This suggests that SafeFix, using a conditional diffusion model, generates images with more internally coherent distributions—that is, the synthetic samples are more semantically consistent with each other. It also shows that LVLM filtering reliably ensures label correctness, resulting in improved downstream classification accuracy. Together, these two key components—the CDM and the LVLM—underscore the semantic fidelity and effectiveness of our augmentation strategy.
Appendix G Future Work
@@
  • Data–selection vs. data–generation. HiBug2 Chen et al. [2025] focuses on discovering failure slices and then retrieving real images from the web to rebalance those slices instead of generating images. A natural extension is to combine the web–retrieval paradigm with SafeFix’s LVLM verifier—e.g. apply the same binary / attribute QA filter to crawled images before retraining. This may further reduce covariate shift without the risk of generative artifacts.
  • Active–learning for efficient selection. Rather than adding a fixed number of images per slice, we can adapt recent active–learning strategies to query either the diffusion model or a search engine for the most informative examples, minimizing GPU time and annotation cost.

### v2

@@ G.4 Evaluating Synthetic Data as a Standalone Training Set
1,000 63.81 81.95
5,000 76.44 90.54
10,000 83.03 90.87
Table 11: Test accuracy (%) on NN synthetic images, trained from scratch. The NN images were split into an 8:1:1 train–val–test partition for this experiment.
In our main experiments, we added $N$ validated synthetic images to $\mathcal{D}_{\mathrm{train}}$ while keeping $\mathcal{D}_{\mathrm{val}}$ and $\mathcal{D}_{\mathrm{test}}$ fixed. To further examine the end-to-end value of our augmentation, we instead randomly split the $N$ synthetic images into an 8:1:1 train–val–test partition and trained ResNet from scratch. As shown in Table 11, our method consistently outperforms HiBug by large margins across all settings. This suggests that SafeFix, using a conditional diffusion model, generates images with more internally coherent distributions—that is, the synthetic samples are more semantically consistent with each other. It also shows that LVLM filtering reliably ensures label correctness, resulting in improved downstream classification accuracy. Together, these two key components—the CDM and the LVLM—underscore the semantic fidelity and effectiveness of our augmentation strategy.
G.5 Threshold settings for identifying rare-case slices
Settings for the rare threshold ($\rho$) and the accuracy-difference threshold ($\epsilon$) are dataset-dependent. As shown in Table 12, with $\rho=0.05$, the number of identified bugs drops from 10 to 4 on CelebA and from 13 to 7 on ImageNet10. For CelebA, increasing $\rho$ to 0.06, 0.07, and 0.08 yields the same result, which raises the pre-fix bug count to 12, while the post-fix count remains stable at 4. Extreme values of $\rho$ become invalid: for ImageNet10, which contains 15 color categories, a value such as $\rho=0.2$ exceeds the $1/15$ ratio of each color and thus no longer reflects a rare case.
Dataset $\rho$ Pre Post
CelebA 0.01 3 2
CelebA 0.05 10 4
CelebA 0.06 12 4
CelebA 0.07 12 4
CelebA 0.08 12 4
CelebA 0.14 15 9
ImageNet10 0.01 4 3
ImageNet10 0.05 13 7
ImageNet10 0.066 16 9
Table 12: Effect of rare threshold $\rho$ on the number of identified bugs before and after repair.
The accuracy-difference threshold $\epsilon$ follows a similar trend. Small values isolate a limited set of failure slices, while moderate values increase the number of identified bugs without affecting the repaired result. Table 13 shows that on both datasets, using $\epsilon$ equals to 0.01, 00.02 and 0.03 yields stable post-fix counts. Extreme thresholds such as $\epsilon=0.1$ become unreliable because they classify many non-critical cases as failures, inflating both pre- and post-fix bug counts.
Dataset $\epsilon$ Pre Post
CelebA 0.01 6 3
CelebA 0.02 8 3
CelebA 0.03 10 4
CelebA 0.10 13 9
ImageNet10 0.01 7 4
ImageNet10 0.02 9 5
ImageNet10 0.03 13 7
ImageNet10 0.10 15 12
Table 13: Effect of accuracy-difference threshold $\epsilon$ on the number of identified bugs before and after repair.
Appendix H Future Work
@@
  • Data–selection vs. data–generation. HiBug2 [chen2025hibug2] focuses on discovering failure slices and then retrieving real images from the web to rebalance those slices instead of generating images. A natural extension is to combine the web–retrieval paradigm with SafeFix’s LVLM verifier—e.g. apply the same binary / attribute QA filter to crawled images before retraining. This may further reduce covariate shift without the risk of generative artifacts.
@@
  • Beyond discrete attributes. Discrete vocabularies of attributes are common failure sources, hence our focus. Future work can extend this direction by handling continuous or fine-grained attributes, supporting multi-attribute interactions, and studying how these repair signals transfer across datasets and model families.